• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 16:03:40 on November16,2000

    Entry number 52207

    keyword=Shift Summary, Nov 16, 2000, 0800-1600

    Shift Summary, Nov 16, 2000, 0800-1600
    ___________________________________

    0800: Run 1350 Continuing, Q^2 = 5.6 (GeV/c)^2 Continuing.
    0816: Run 1350 ended. 1.0M events
    Run 1351 started.
    0831: Run 1351 ended. 0.223M event.
    MCC takes away beam for spot change for Hall C.
    -- Put in 15cm dummy target.
    0956: Run 1352 started. Dummy 15 cm target.
    1045: Run 1352 ended. 0.314M events
    -- Put in LH2 target
    1057: Run 1353 started.
    1152: Run 1353 ended. 1M events
    Run 1354: junk run. Channel 8.1.16 FPP HV tripped.
    1158: Run 1355 started.
    1303: Run 1355 ended. 1M events
    Run 1356 started
    1402: Run 1356 ended. 1M events
    Run 1357 started
    1445: Mcc takes away beam to fix problem. Estimate ~ 5 mins.
    1520: Run 1357 ended. 1M events
    Run 1358 started
    1600: Run 1358 continuing (0.75M events and counting)